OpenStudy (anonymous):

Find the exact value of arccos (sin (pi/6)). For full credit, explain your reasoning.

OpenStudy (mathteacher1729):

http://en.wikipedia.org/wiki/File:Unit_circle_angles_color.svg If you can print and laminate that, go for it. It's very useful. Also this: http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

if i understood the question then sin(pi/6)=1/2, arccos(1/2)=pi/3

OpenStudy (anonymous):

@mathsmind so the final answer will be pi/3 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think everyone here knows the answer but they are trying to teach you

OpenStudy (anonymous):

how should i explain the reasoning? yeahh

OpenStudy (anonymous):

well as the guys said you can look at the triangle, from their you find the value of sin(pi/6). then the arccos would be the value obtained from the triangle and unit circle...

OpenStudy (anonymous):

i don't know if you know arccos triangle as well, which needs a bit of explination...

OpenStudy (anonymous):

let me explain this in an algabreac way,,, if theta =arccos(x) and x=sin(pi/6) then all this is saying you are dealing with the same triangle not different one by definition.. I hope this makes sense...

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):



OpenStudy (anonymous):